OpenStudy (anonymous):

1. simplified form of x^4-81/x+3 2. simplified form of (x^2yz)^2(xy^2z^2)/(xyz)^2

OpenStudy (lgbasallote):

1) hmm if you notice you have difference of two squares there... (x^2)^2 - (9)^2 yes? try factoring that out..

OpenStudy (anonymous):

\[\Large \frac{(x^2yz)^2(xy^2z^2)}{(xyz)^2}\] Use the rule.. \[\LARGE \frac{a^m}{a^n}=a^{m-n}\] And you'll have... \[\LARGE y^2z^2 x^3\]

OpenStudy (anonymous):

wait im confused about the first one

OpenStudy (anonymous):

so would the first one be (x^2 +9)(x^2-9)??

OpenStudy (anonymous):

@haileystowers so you don't understand the first question??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\LARGE x^4-\frac{81}{x}+3\] .. hmmm actually , I don't know what to simplify there, do you have multiple choices ?? :S

OpenStudy (anonymous):

@lgbasallote can you explain please how do you simplify this one, It seems I don't get it :( ... (first question ! )

OpenStudy (lgbasallote):

it's (x^4 - 81)/(x+3) =))) now..doesnt that make more sense /:)

OpenStudy (lgbasallote):

you finish it @Kreshnik ^_^

OpenStudy (anonymous):

auff... I thought it was the one I wrote a few minutes ago LOL ... thnx... here you go ... \[\LARGE \frac{x^4-81}{x+3}=\frac{(x^2)^2-9^2}{x+3}\] now use this formula... \[\LARGE a^2-b^2=(a-b)(a+b)\] \[\LARGE \frac{(x^2-9)(x^2+9)}{x+3}\] Final Answer !!